OpenStudy (anonymous):

True or False, If x>0, then (ln x)^6 = 6 ln x

jimthompson5910 (jim_thompson5910):

Hint: Let x = e

OpenStudy (anonymous):

false... but if ln(x^6) = 6 lnx

OpenStudy (anonymous):

Its true isnt it?

jimthompson5910 (jim_thompson5910):

(ln x)^6 = 6*ln x (ln e)^6 = 6*ln e (1)^6 = 6*1 1 = 6 So it's false

OpenStudy (anonymous):

thanks! I get it now! :)

jimthompson5910 (jim_thompson5910):

np